# Twisted Cherednik systems and non-symmetric Macdonald polynomials

### v1

@@ 2.1 A toy example
Symmetric eigenfunctions are labeled by the Young diagrams (ordered partitions) $\lambda^{+}$, i.e. ordered sequences of positive integers $\lambda_{1}\geq\lambda_{2}\geq\ldots>0$, describing partitions of the level $|\lambda|:=\sum_{i}\lambda_{i}$:
@@ 2.2.1 Iterative construction of non-symmetric polynomials
where the Schur polynomial $S_{R}$ at the r.h.s. with the partition R=[k1]R=[k-1] depends only on two $x$-variables and can not be symmetric in all the $n+1$.
@@ 3.2 Orthogonality relations
where $\lambda$ is a partition with $n$ parts (unordered, and some of the parts may be zero). If there are two partitions, $\lambda$ and $\mu$, $\lambda>\mu$ if the ordered partition $\lambda^{+}>\mu^{+}$ (e.g., in accordance with the lexicographic order), and if the ordered partitions coincide, one compares the minimal length of permutations of the symmetric group ${\cal S}_{n}$ that allow one to make an ordered partition. The less is the length, the larger is partition. In other words, the largest one is $\lambda^{+}$, and in the sum in (26) all $\mu$ with $\mu^{+}=\lambda^{+}$ are present. The next smaller one is any one $\mu_{i}$ from the set of $\{\mu_{i}=\sigma_{i}(\mu^{+})\}$, $i=1,\ldots,n-1$ given by a single elementary transposition. Such $\mu_{i}$ does not include into the sum only $\mu^{+}$ and all the elements of this set, etc. This is called Bruhat order.
@@ 3.3 Non-symmetric Macdonald polynomials as eigenfunctions of the Cherednik operators
where $\Lambda_{\lambda}^{(i)}$ are eigenvalues. If one considers solutions of a given homogeneity $p$ in $x_{i}$, these equations have the number of non-trivial solutions as many as the number of unordered partitions $\lambda$ of $p$ in $n$ parts, which are just the non-symmetric Macdonald polynomials $E_{\lambda}$. Note that, with the notation used here, the polynomials are obtained with opposite numeration of $x_{i}$ as compared with [24]:
@@
$E_{[0,2,0]}$ $=$ $x_{2}^{2}+{q^{2}t(1-t)\over(1-q^{2}t^{2})}x_{3}^{2}+{1-t\over 1-qt}x_{1}x_{2}+{t(1-t)^{2}\over(1+qt)(1-qt)^{2}}x_{1}x_{3}+{q(1-t)(1+qt-qt^{2}-q^{2}t^{2})\over(1+qt)(1-qt)^{2}}x_{2}x_{3}$
@@ 3.5 Creation operators
Note that operators $T_{i}$ allows one to construct the non-symmetric Macdonald polynomials recursively. Indeed, the action of this operators just permutes the $i$-th and $(i+1)$-th parts of the partition so that
@@
So far, we had operators that permuted parts of the partition $\lambda$. Now we construct the operator that increases partitions:
@@ 5.1 Eigenfunctions of higher Cherednik Hamiltonians
are again labeled by unordered partitions $\lambda$. Moreover, the solutions turns out to be proportional to the non-symmetric Macdonald polynomials:
@@ 5.3 Constructing polynomial eigenfunctions
and $|\lambda|$ even, and realize that solutions are again numbered by unordered partitions!
@@
and $|\lambda|$ odd to realize that solutions are also numbered by unordered partitions! It completes the construction of 2-twisted non-symmetric Macdonald polynomials $E_{\lambda}^{(2)}(x_{1},x_{2})$.
@@ 6.1 General solution at $n=2$
At $n=2$ and higher $a$, the whole construction becomes different, and reduces to the construction of $p=2$ case described in the previous section in a peculiar way. The general polynomial solution of the eigenfunction equations at $t=q^{-m}$ is again based on the lowest grade solution $\Omega^{(a)}(m;x_{1},x_{2})$, which is
@@ 6.3 Extension to higher $n$
At level $\lambda=1$, there are three polynomial solutions associated with three possible partitions [0,0,1], [0,1,0] and [1,0,0]. They have the form
@@
where we use the notation $\{x\}:=1-x$. Note that these solutions are still polynomials, though it is not evident at all: this is a peculiar property of the Baker-Akhiezer function $\Omega^{(a)}$.
Similarly, at level $|\lambda|=2$:
@@
$\Xi^{(a)}_{\lambda}:=\left(\prod_{i=1}x_{i}^{\lambda_{i}\over a}q^{\lambda_{i}(\lambda_{i}-1)\over 2a}\right)\ \Omega^{(a)}(q,t;\{q^{\lambda_{i}}x_{i}\})=\prod_{i=1}\left(x_{i}^{1\over a}q^{\hat{D}_{i}}\right)^{\lambda_{i}}\ \Omega^{(a)}(q,t;\{q^{\lambda_{i}}x_{i}\})$ (114)
@@
Then, one can naturally expect the general formula to be of the form (notice the triangular structure)
$\boxed{\psi_{\lambda}=\sum_{\mu\leq\lambda}c_{\lambda\mu}\left(\prod_{i,j}{\Big\{{a_{ij}(\mu)x_{i}\over x_{j}}\Big\}\over\Big\{{b_{ij}(\mu)x_{i}\over x_{j}}\Big\}}\right)\Xi^{(a)}_{\mu}}$ (116)
where $a_{ij}(\mu)$, $b_{ij}(\mu)$ are monomials of $q$ and $t$. All these quantities $c_{\lambda\mu}$, $a_{ij}(\mu)$ and $b_{ij}(\mu)$ do not depend on $a$. Moreover, at $t=q^{-m}$, the coefficients $c_{\lambda\mu}$ are just ratios of $q$-numbers, and $\psi_{\lambda}$ is a polynomial (!), while individual terms in the sum at each $\mu$ are not.
@@
Denote the number of solutions of the eigenvalue equations through $p(\lambda,n)$. We checked at various values of $a$, $n$ and $\lambda$ that $p(\lambda,n)$ does not depend on $a$, and the generating function of these numbers is
$\sum_{\lambda=0}p(\lambda,n)z^{\lambda}={1-z^{n}\over(1-z)^{n}}$ (117)
In particular, $p(\lambda,n)$ exactly matches the number of “basic” non-symmetric Macdonald polynomials, which are defined either as those not having a prefactor $\prod_{i=1}^{n}x_{i}^{\alpha}$, or those labeled by partitions containing, at least one zero part. Taking into account this fact and presumable formula (116),
@@
We checked this conjecture at various particular values of $a$ and $n$, it perfectly works.
@@ 7.1 Summary
  • The generic eigenfunctions can be realized (116) as linear sums of $\Omega^{(a)}$ (multiplied by proper monomials made of $x_{i}^{1\over a}$ and $q^{1\over a}$) with expansion functions (rational functions) that do not depend on $a$. Hence, only the ground state functions $\Omega^{(a)}$ control a peculiar twisting. This universality reflects an $SL(2,\mathbb{Z})$ symmetry of the DIM (automorphism Miki [30]) and Cherednik algebras.
@@ 7.2 Discussion
  • The eigenfunctions of the twisted Cherednik system are conjectured to be described by formula (116). However, the parameters $a_{ij}(\mu)$, $b_{ij}(\mu)$ yet to be determined in order to achieve at arbitrary $n$ the concreteness similar to formula (104) in the case of $n=2$.

### v2

@@ 2.1 A toy example
Symmetric eigenfunctions are labeled by the Young diagrams (ordered partitions) $\lambda^{+}$, i.e. ordered sequences of positive integers $\lambda_{1}\geq\lambda_{2}\geq\ldots>0$, describing weak compositionsof the level $|\lambda|:=\sum_{i}\lambda_{i}$:
@@ 2.2.1 Iterative construction of non-symmetric polynomials
where the Schur polynomial $S_{R}$ at the r.h.s. with the Young diagram R=[k1]R=[k-1] depends only on two $x$-variables and can not be symmetric in all the $n+1$.
@@ 3.2 Orthogonality relations
where $\lambda$ is a weak compositionwith $n$ parts (unordered, and some of the parts may be zero). If there are two weak compositions, $\lambda$ and $\mu$, $\lambda>\mu$ if the ordered partition $\lambda^{+}>\mu^{+}$ (e.g., in accordance with the lexicographic order), and if the ordered partitions coincide, one compares the minimal length of permutations of the symmetric group ${\cal S}_{n}$ that allow one to make an ordered partition. The less is the length, the larger is weak composition. In other words, the largest one is $\lambda^{+}$, and in the sum in (26) all $\mu$ with $\mu^{+}=\lambda^{+}$ are present. The next smaller one is any one $\mu_{i}$ from the set of $\{\mu_{i}=\sigma_{i}(\mu^{+})\}$, $i=1,\ldots,n-1$ given by a single elementary transposition. Such $\mu_{i}$ does not include into the sum only $\mu^{+}$ and all the elements of this set, etc. This is called Bruhat order.
@@ 3.3 Non-symmetric Macdonald polynomials as eigenfunctions of the Cherednik operators
[…]
@@
$E_{[0,2,0]}$ $=$ $x_{2}^{2}+{q^{2}t(1-t)\over(1-q^{2}t^{2})}x_{3}^{2}+{1-t\over 1-qt}x_{1}x_{2}+{q^{2}t(1-t)^{2}\over(1+qt)(1-qt)^{2}}x_{1}x_{3}+{q(1-t)(1+qt-qt^{2}-q^{2}t^{2})\over(1+qt)(1-qt)^{2}}x_{2}x_{3}$
@@ 3.5 Creation operators
Note that operators $T_{i}$ allows one to construct the non-symmetric Macdonald polynomials recursively. Indeed, the action of this operators just permutes the $i$-th and $(i+1)$-th parts of the weak compositionso that
@@
[…]
@@ 5.1 Eigenfunctions of higher Cherednik Hamiltonians
[…]
@@ 5.3 Constructing polynomial eigenfunctions
and $|\lambda|$ even, and realize that solutions are again numbered by weak compositions!
@@
and $|\lambda|$ odd to realize that solutions are also numbered by weak compositions! It completes the construction of 2-twisted non-symmetric Macdonald polynomials $E_{\lambda}^{(2)}(x_{1},x_{2})$.
@@ 6.1 General solution at $n=2$
At $n=2$ and higher $a$, the whole construction becomes different, and reduces to the construction of $a=2$ case described in the previous section in a peculiar way. The general polynomial solution of the eigenfunction equations at $t=q^{-m}$ is again based on the lowest grade solution $\Omega^{(a)}(m;x_{1},x_{2})$, which is
@@ 6.3 Extension to higher $n$
At level $|\lambda|=1$, there are three eigenfunctions associated with three possible weak compositions [0,0,1], [0,1,0] and [1,0,0]. They have the form
@@
where we use the notation $\{x\}:=1-x$. Note that these solutions are still polynomials at $t=q^{-m}$, though it is not evident at all: this is a peculiar property of the Baker-Akhiezer function $\Omega^{(a)}$, which deserves further studying. One can see that the coefficients in each of these expressions contain the same number of fractions, and this number in $\psi_{\lambda}^{(a)}$ is determined by the minimal number of permutations needed to obtain $\lambda$ from $w_{0}\lambda$, where $w_{0}$ is the longest permutation in permutation group ${\cal S}_{n}$. For instance, in $\psi_{[0,0,1]}^{(a)}$ there are no fractions, since $0,0,1]=w_{0}[1,0,0]$. In $\psi_{[0,1,0]}^{(a)}$ there is one fraction, and [0,1,0] is obtained from [0,0,1] by one permutation minimally, while in $\psi_{[1,0,0]}^{(a)}$ there are two fractions, and [1,0,0] is obtained from [0,0,1] minimally by two permutations.
Similarly, at level $|\lambda|=2$, the simplest eigenfunctions are
@@
$\Xi^{(a)}_{\lambda}:=\left(\prod_{i=1}x_{i}^{\lambda_{i}\over a}q^{\lambda_{i}(\lambda_{i}-1)\over 2a}\right)\ \Omega^{(a)}(q,t;\{q^{\lambda_{i}}x_{i}\})=\prod_{i=1}\left(x_{i}^{1\over a}q^{\hat{D}_{i}}\right)^{\lambda_{i}}\ \Omega^{(a)}(q,t;\{x_{i}\})$ (114)
Thus defined quantity automatically takes into account that multiplying a solution with a factor of $\prod_{i=1}^{n}x_{i}^{\alpha}$ gives rise to another solution, and shifts all entries in the weak composition by $\alpha$: $\lambda_{i}\to\lambda_{i}+\alpha$. This follows from the property $\Omega^{(a)}(\{q^{\alpha}x_{i}\})\sim\Omega^{(a)}(\{x_{i}\})$. Thus, now one can immediately deal with all possible weak compositions, not obligatory with those having at least one zero part.
@@
[…]
$\psi_{[0,2,0]}^{(a)}$ $=$ ${\Big\{{qtx_{2}\over x_{3}}\Big\}\over\Big\{{qx_{2}\over x_{3}}\Big\}}{\Big\{{qtx_{2}\over x_{1}}\Big\}\over\Big\{{qx_{2}\over x_{1}}\Big\}}{\Big\{{tx_{2}\over x_{3}}\Big\}\over\Big\{{x_{2}\over x_{3}}\Big\}}\Xi^{(a)}_{[0,2,0]}+{(1-t)\over(1-q^{2}t^{2})}{\Big\{{q^{2}t^{2}x_{3}\over x_{2}}\Big\}\over\Big\{{qx_{3}\over x_{2}}\Big\}}{\Big\{{qtx_{3}\over x_{2}}\Big\}\over\Big\{{x_{3}\over x_{2}}\Big\}}{\Big\{{qtx_{3}\over x_{1}}\Big\}\over\Big\{{qx_{3}\over x_{1}}\Big\}}\Xi^{(a)}_{[0,0,2]}$+ (116)
$+$ ${q(1-t)\over(1-qt)(1+t)}\left({(1+q)(1-qt^{2})\over(1-q^{2}t^{2})}{\Big\{{qt^{2}x_{3}\over x_{2}}\Big\}\over\Big\{{qx_{3}\over x_{2}}\Big\}}{\Big\{{tx_{2}\over x_{3}}\Big\}\over\Big\{{qx_{2}\over x_{3}}\Big\}}{\Big\{{tx_{2}\over x_{1}}\Big\}\over\Big\{{x_{2}\over x_{1}}\Big\}}+{\Big\{{tx_{2}\over x_{3}}\Big\}\over\Big\{{qx_{2}\over x_{3}}\Big\}}{\Big\{{t^{2}x_{3}\over x_{1}}\Big\}\over\Big\{{x_{3}\over x_{1}}\Big\}}{\Big\{{tx_{1}\over x_{2}}\Big\}\over\Big\{{x_{1}\over x_{2}}\Big\}}\right)\Xi^{(a)}_{[0,1,1]}$+
$+$ ${(1-t)^{2}\over(1-qt)(1-q^{2}t^{2})}{\Big\{{q^{2}t^{2}x_{3}\over x_{2}}\Big\}\over\Big\{{x_{3}\over x_{2}}\Big\}}{\Big\{{tx_{1}\over x_{3}}\Big\}\over\Big\{{x_{1}\over qx_{3}}\Big\}}{\Big\{{tx_{1}\over x_{2}}\Big\}\over\Big\{{x_{1}\over x_{2}}\Big\}}\Xi^{(a)}_{[1,0,1]}+{(1-t)\over(1-qt)}{\Big\{{tx_{2}\over x_{3}}\Big\}\over\Big\{{x_{2}\over x_{3}}\Big\}}{\Big\{{tx_{1}\over x_{3}}\Big\}\over\Big\{{x_{1}\over x_{3}}\Big\}}{\Big\{{tx_{1}\over x_{2}}\Big\}\over\Big\{{x_{1}\over qx_{2}}\Big\}}\Xi^{(a)}_{[1,1,0]}$
At last, the sixth remaining eigenfunction at this level is
[…]
$+$ $\left.{\Big\{{tx_{3}\over x_{2}}\Big\}\over\Big\{{x_{3}\over x_{2}}\Big\}}{\Big\{{qtx_{3}\over x_{2}}\Big\}\over\Big\{{qx_{3}\over x_{2}}\Big\}}{\Big\{{qtx_{3}\over x_{1}}\Big\}\over\Big\{{qx_{3}\over x_{1}}\Big\}}{\Big\{{q^{2}tx_{3}\over x_{1}}\Big\}\over\Big\{{x_{3}\over x_{1}}\Big\}}\Xi^{(a)}_{[0,0,2]}\right)+{q(1+q)(1-t)^{2}\over(1-qt)(1-q^{2}t)}{\Big\{{tx_{3}\over x_{2}}\Big\}\over\Big\{{qx_{3}\over x_{2}}\Big\}}{\Big\{{tx_{2}\over x_{3}}\Big\}\over\Big\{{qx_{2}\over x_{3}}\Big\}}{\Big\{{qtx_{3}\over x_{1}}\Big\}\over\Big\{{x_{3}\over x_{1}}\Big\}}{\Big\{{qtx_{2}\over x_{1}}\Big\}\over\Big\{{x_{2}\over x_{1}}\Big\}}\Xi^{(a)}_{[0,1,1]}+$
$+$ ${q(1+q)(1-t)\over(1-q^{2}t)}\left({\Big\{{tx_{3}\over x_{2}}\Big\}\over\Big\{{x_{3}\over x_{2}}\Big\}}{\Big\{{tx_{1}\over x_{3}}\Big\}\over\Big\{{qx_{1}\over x_{3}}\Big\}}{\Big\{{qtx_{3}\over x_{1}}\Big\}\over\Big\{{qx_{3}\over x_{1}}\Big\}}{\Big\{{tx_{1}\over x_{2}}\Big\}\over\Big\{{x_{1}\over x_{2}}\Big\}}\Xi^{(a)}_{[1,0,1]}+{\Big\{{tx_{2}\over x_{3}}\Big\}\over\Big\{{x_{2}\over x_{3}}\Big\}}{\Big\{{tx_{1}\over x_{2}}\Big\}\over\Big\{{qx_{1}\over x_{2}}\Big\}}{\Big\{{qtx_{2}\over x_{1}}\Big\}\over\Big\{{qx_{2}\over x_{1}}\Big\}}{\Big\{{tx_{1}\over x_{3}}\Big\}\over\Big\{{x_{1}\over x_{3}}\Big\}}\Xi^{(a)}_{[1,1,0]}\right)$
Note that the pattern with the number of fraction in each term persists: in $\psi_{[0,1,1]}^{(a)}$ there are no fractions, in $\psi_{[1,0,1]}^{(a)}$ there is one ([1,0,1] is obtained from [0,0,1] minimally by one permutation), in $\psi_{[1,1,0]}^{(a)}$ there are two fractions ([1,1,0] is obtained from [0,0,1] minimally by two permutation). Similarly, as soon as in $\psi_{[0,0,2]}^{(a)}$ there are two fractions, in $\psi_{[0,2,0]}^{(a)}$ there are three, and in $\psi_{[2,0,0]}^{(a)}$ there are four fractions.
@@
The same structure of eigenfunctions emerges at higher $n$ and $\lambda$. Thus, one can naturally expect the general formula for the eigenfunction to be of the form (notice the triangular structure)
$\boxed{\psi_{\lambda}=\sum_{\mu\leq\lambda}F_{\lambda\mu}(x)\ \Xi^{(a)}_{\mu}}$ (118)
and all $F_{\lambda\mu}(x)$’s do not depend on $a$.
Each $F_{\lambda\mu}(x)$ is a homogeneous rational function of $x_{i}$’s, which generally is a sum of products of the form
$F_{\lambda\mu}(x)\sim\sum\prod_{(i,j)}^{N_{\lambda}}{\Big\{{a_{ij}tx_{i}\over x_{j}}\Big\}\over\Big\{{b_{ij}x_{i}\over x_{j}}\Big\}}$ (119)
and the number of fractions $N_{\lambda}$ in these products is the same for all $\mu$ at fixed $\lambda$, and $N_{\lambda^{+}}-N_{\lambda}$ is equal to the minimal length of permutation that brings the weak composition $\lambda$ to $\lambda^{+}$. Here $a_{ij}$, $b_{ij}$ are monomials of $q$ and $t$. Moreover, at $\bf t=q^{-m}$}, the coefficients in front of products (119) entering $F_{\lambda\mu}(x)$ are ratios of $q$-numbers [^8] From this, it is clear from the very beginning that $F_{[0,0,2][0,1,1]}(x)$ in (116) contains two terms: the coefficient in front of $x_{2}x_{3}$ in (3.3) is not a ratio of $q$-numbers because of the factor $(1+qt-qt^{2}-q^{2}t^{2})$, which is a sum of two $q$-numbers, and implies a non-trivial multiplicity in this case. (up to possible monomials of $q$), and $\psi_{\lambda}$ is a polynomial (!), while individual terms in the sum at each $\mu$ are not.
@@
Denote the number of solutions of the eigenvalue equations through $p(\lambda,n)$. We checked at $t=q^{-m}$ and at various values of $a$, $n$, $\lambda$ and $m>|\lambda|$ that $p(\lambda,n)$ does not depend on $a$ and $m$. Taking into account this fact and presumable formula (118),
@@
We tested this conjecture at various particular values of $a$ and $n$, it perfectly works.
@@ 7.1 Summary
  • The generic eigenfunctions can be realized (118) as linear sums of $\Omega^{(a)}$ (multiplied by proper monomials made of $x_{i}^{1\over a}$ and $q^{1\over a}$) with expansion functions (rational functions of $x_{i}$) that do not depend on $a$. Hence, only the ground state functions $\Omega^{(a)}$ control a peculiar twisting. This universality reflects an $SL(2,\mathbb{Z})$ symmetry of the DIM (automorphism Miki [30]) and Cherednik algebras.
@@ 7.2 Discussion
  • The eigenfunctions of the twisted Cherednik system are conjectured to be described by formula (118). However, the explicit form of the rational functions $F_{\lambda\mu}(x)$ yet to be further specified in order to achieve at arbitrary $n$ the concreteness similar to formula (104) in the case of $n=2$.